OpenStudy (anonymous):

Write the slope -intercept equation for the line that passes through (-8,-11) and is perpendicular to -2x-7y=-14?

OpenStudy (anonymous):

does anyone know this one??

OpenStudy (anonymous):

the line perpendicular to -2x-7y=-14 is given by 7x-2y+k=0. as the line passes through (-8,-11) ,substitute it in the above equation. we get..... -56+22+k=0 k=34 so the required equation of the line is 7x-2y+34=0.